ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35J, 45E Potthast, Roland, University of Göttingen On mixed reciprocity relations in inverse scattering. Reciprocity relations are of great importance in the treatment of direct and inverse scattering problems. Mixed reciprocity relations relate the far field pattern of incident point-sources with the scattered field for incident plane waves. Here we derive mixed reciprocity relations for some acoustic and electromagnetic scattering problems and show their applicability to solve inverse obstacle scattering problems.